in the 1920 the record for a certain race was 46.3 sec. in 1930 it was 46.1sec. let R(t)= the record in the race and t= the number of years since 1920 .

a)find a linear function that finds the data
b)use the function in a) to predict the record in 2003 and 2006
c)find the year when the record will be 44.50

Answer by graphmatics(170)

The decrease in sec of race speed reduction of t years is,
f(t) = (46.3-46.1)*(t/10)
f(t) = 0.2*(t/10)
f(t) = 0.2*t/10
f(t) = 0.2*t/10
___________________________________________________
so for 2003
f(2003-1930) = 0.2*((2003-1930)/10))
f(t) = 1.46 secs will be the decrease, so the record speed is
46.3 - 1.46 = 44.84
___________________________________________________
To find the year when the record will be 44.50
46.3 - f(t-1930) = 44.50
-f(t-1930) = 44.50 - 46.3
-f(t-1930) = -1.8
f(t-1939) = 1.8
0.2*(t-1930)/10 = 1.8
0.2*t = 18 + 0.2*1930
t = (18 + 386)/0.2
t = 2020